Obtain the differential equation by eliminating arbitrary constant.





A ln y = Be^x²

1
Expert's answer
2021-11-09T16:48:53-0500

"A ln y = Be^{x^2} ...(1)\\"

Differentiating both sides w.r.t "x" , we get

"\frac{A}{y}.y'=2xBe^{x^2}\\
\frac{A}{y}.y'=2x.A ln y\ [\text{From eqn. (1)}]\\
y'=2xy.lny\\
y'-2xy.lny=0"